# Distributional Random Forests for Complex Survey Designs on Reproducing Kernel Hilbert Spaces

### v1

@@ 3.2 Splitting Rule
Fix a node $P_{a}$ and a candidate split $\theta$ that induces children $C_{L},C_{R}$ with $N_{P_{a}}:=|\{i\leq N:X_{i}\in P_{a}\}|$, $N_{\ell}:=|\{i\leq N:X_{i}\in C_{\ell}\}|$. Suppose there exists $\alpha\in(0,1/2]$ such that $\frac{N_{L}}{N_{P_{a}}},\ \frac{N_{R}}{N_{P_{a}}}\in[\alpha,1-\alpha].$ Let $M_{N}(\theta)$ be the finite-population split score, $M_{n_{s}}(\theta)$ its design-weighted sample analogue, and $M_{n_{s}}^{*(M)}(\theta)$ the resampled score computed using averaged multipliers $\bar{n}_{i}^{*}:=M^{-1}\sum_{b=1}^{M}n_{b,i}^{*}$ (equivalently $\bar{\delta}_{i}:=\bar{n}_{i}^{*}-1$). Under (D1)–(D4), (R1)–(R3), and (K1)–(K2), if the multiplier increments $\delta_{b,i}:=n_{b,i}^{*}-1$ are conditionally sub-gamma and negatively dependent given $\xi$, which are satisfied by the methods described in Section A.5.1, then
$\bigl|M_{n_{s}}^{*(M)}(\theta)-M_{N}(\theta)\bigr|=O_{p}\!\left(N_{P_{a}}^{-1/2}\right)\;+\;O_{p}\!\left(\frac{1}{\sqrt{M}}\sum_{\ell\in\{L,R\}}n_{\mathrm{eff},\ell}^{-3/2}\right),$
where $n_{\mathrm{eff},\ell}$ is the design-weighted effective sample size ([49]) in child $\ell$ induced by $\theta$.
@@
Proposition 3.3 separates two sources of split-score fluctuation: an intrinsic design term $|M_{n_{s}}(\theta)-M_{N}(\theta)|=O_{p}(N_{P_{a}}^{-1/2})$ which concerns the sampling noise already present even with fixed $n^{*}$, and a resampling term $|M_{n_{s}}^{*(M)}(\theta)-M_{n_{s}}(\theta)|=O_{p}\!\big(M^{-1/2}\sum_{\ell\in\{L,R\}}n_{\mathrm{eff},\ell}^{-3/2}\big)$.
The cross-over threshold occurs when these terms balance: if $M\gtrsim\tilde{M}$ where $\tilde{M}\asymp N_{P_{a}}\!\big(\sum_{\ell\in\{L,R\}}\\ n_{\mathrm{eff},\ell}^{-3/2}\big)^{2},$ then the design term dominates and increasing $M$ further yields negligible stability gains. At deep nodes (small $n_{\mathrm{eff},\ell}$), this threshold is reached quickly: for instance, if both children have $n_{\mathrm{eff},\ell}\approx 100$ and $N_{P_{a}}\approx 200$, then $\tilde{M}\approx 8$ suffices. Beyond this point, split stability is governed by the effective sample sizes—controlled via minimum-node-size constraints, shallower depths, or larger subsampling fractions qq—rather than by the number of resamples.
@@ 3.3 Prediction and Aggregation
    There exist constants $0<m<M<\infty$ such that for all radius r, of some $L(\mathbf{x})$ for the given $\mathbf{x}$ such that $0<r\leq k_{N}$, $mr^{p}\leq\#\{i:X_{i}\in B_{r}(\mathbf{x})\}/N\leq Mr^{p}$, where $B_{r}(\mathbf{x})$ the ball centered at $\mathbf{x}$ with radius $r$.
@@
    For any $L^{*}$ found by $D_{split}$, when populating the values with $D_{est}$, all leaves receive at least $n_{L^{*},min}$ subjects. In other words, $n_{L^{*}(\mathbf{x})}=\sum_{i}^{N}\xi_{i}I(X_{i}\in L^{*}(\mathbf{x}))\geq n_{L^{*},min}>0$.
@@
    For any $\mathbf{x}$ and some deterministic chosen sequence $k_{N}\to 0$ yet $k^{p}_{N}N\to\infty$ as $N\to\infty$, there exists $t_{N}\to 0$ such that $P_{L}(diam(L(\mathbf{x}))\geq k_{N})\leq t_{N}$, where $iam(L(\mathbf{x}))=\sup_{x_{1},x_{2}\in L(\mathbf{x})}d(x_{1},x_{2}), dd$ some metric on $\mathcal{X}^{p}$.
@@
    $\sup_{x\in\mathcal{X}_{1}\cup\mathcal{X}_{2}}\mathbb{P}\{X\in L^{*}(x)\,\Delta\,L(x)\}=o_{p}(d_{N})$ for some $d_{N}=o(k_{N}^{p})$, where $\Delta$ denotes symmetric difference: $A\Delta B:=(A\setminus B)\;\cup\;(B\setminus A)$.
Denote $\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ the estimate from SDRF. Then $d_{\mathbf{k}}(\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}),{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}))\to_{p_{d}}0$ as $N,B\to\infty$.
@@
Regularization–type conditions (A1)–(A3) can be controlled via adaptive stopping criteria. (A1) imposes a geometric control on the shape of $L$ so that they do not be stretched. (A2) requires a local mass control for the local empirical density to be well-behaved. (A3) ensures the local density estimate stays bounded. (B1)-(B3) formalize how the partitions produced by the algorithm behave as the finite population size grows. (B1) is a shrinking diameter assumption that mirrors the adaptive concentration result proved for i.i.d. forests ([94]). (B2) requires that the thickness of the error band be negligible relative to the leaf’s own diameter. In other words, the rate at which the leaf $L^{*}(\cdot)$ function estimated from the bootstrap sample approaches the leaf $L(\cdot)$ estimated if using the full finite population must be faster than the oracle leaf’s own shrinking-volume rate. Under the required relationship between $d_{N}$ and $k_{N}$, together with (A1) regularity conditions on the geometry, the symmetric-difference region $L^{*}(x)\triangle L(x)$ has measure $O_{p}(d_{N}/k_{N}^{p})$, which implies that both the tree estimator $\widehat{T}^{*}_{N}(L_{\xi}^{*}(x))$ and its conditional variance coincide with their oracle counterparts up to $o_{p}(N^{-1/2})$ and $o_{p}(N^{-1})$, respectively. Among all the assumptions, (B2) is the hardest to directly control. Thus we provided Theorem 3.2 that partially shed light on conditions under which (B2) can be reasonably assumed to hold.
@@
Suppose $\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{N}$ i.i.d. realizations from $P_{\Omega}(\cdot)$. Suppose $\mathbf{k}$ satisfying (K1)–(K3), and $\mathcal{Y}$ locally compact Polish (satisfying (S1)). Denote ${P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ the true conditional measure of $Y|X=\mathbf{x}$ in the design space and $\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ an estimate for it. Denote $P_{\omega}(\mathbf{x})$ the true conditional measure of $Y|X=\mathbf{x}$ in the super population space. Then for any specified $\mathbf{x}\in\mathcal{X}$ and $\Psi:(\mathcal{M}_{b}(\mathcal{Y}),\;d_{\mathbf{k}})\to(\mathcal{V},d)$ a measurable functional continuous at $P^{N}(Y|X=\mathbf{x})$ for a.e. $\omega\in\Omega$, if   $d_{\mathbf{k}}(\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}),{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}))\to_{p_{d}}0,\text{ then }$
@@
Corollary 3.5 reduces design– and model–based consistency of $\widehat{\Psi}_{N}(\mathbf{x})$ to control of $\widehat{P}_{S^{N}\mid\omega}(\mathbf{x})$ in the $d_{\mathbf{k}}$-metric, together with continuity of the functional $\Psi$ at the relevant target law. Many standard summaries satisfy this requirement under mild regularity. For example, when $\mathcal{Y}=\mathbb{R}^{d}$, linear functionals such as conditional means $\Psi(P)=\int y\,dP(y)$ and, more generally, $\Psi_{g}(P)=\int g(y)\,dP(y)$ for $g\in C_{b}(\mathcal{Y})$ are continuous under weak convergence. For distributional functionals, $\Psi_{t}(P)=F_{P}(t)$ and $\Psi_{t}(P)=P(Y>t)$ are continuous at $P$ whenever $P(\{t\})=0$. The quantile functional $\Psi_{q}(P)=F_{P}^{-1}(q)$ is continuous at $P$ whenever the $q$-quantile is unique. Equivalently, that $F_{P}$ has no flat segment at level $q$. The kernel mean element itself and its norm $\Psi(P)=\|\mu_{\mathbf{k}}(P)\|_{\mathcal{H}}$, and the MMD between the true conditional $P$ to a reference distribution $Q$, $d_{\mathbf{k}}(P,Q)$, are also examples.
@@ 3.4 Implementation Considerations
  • For the re–sampling, ([22]) showed that asymptotically, generating one pseudo population and then drawing $B$ samples from it (equivalently, realizing $B$ times $n^{*}$ conditioning on the realized pseudo population) is asymptotically equivalent to realizing independently $B$ pseudo population “and” the re-sample.
  • RBF–type kernel is a common class of kernels that is translation–invariant. Median-heuristic, i.e., specifying the bandwidth $\sigma>0$ to be the median of pairwise sampled distance, is a commonly adopted, and reasonable choice ([35, 64]).
  • Hyperparameters governing the stopping criteria are introduced to help ensure regularity-type conditions (A1)–(A3). For finite samples, it may be preferable to use a moderate number of shallow trees to keep local sample weights well-controlled. Moreover, when the number of PSUs is large but PSU sizes are small, one may consider omitting second-stage resampling (Section A.5.1).
  • Because SDRF yields a kernel-type estimator, one could further post-process $\widehat{P}(Y\mid X=\mathbf{x})$ (or $\widehat{\Psi}(\mathbf{x})$) with an additional kernel-smoothing step [92] in $\mathbf{x}$ to reduce finite-sample variability and stabilize the resulting estimates, particularly when it is reasonable to assume that the underlying target varies smoothly in $\mathbf{x}$.
To conclude, up to this point, we have illustrated the necessity of an appropriate method when data arise under sampling designs as opposed to being i.i.d.; described in detail our proposed algorithm as a design-aware solution; proved both types of consistency for the resulting estimator under reasonable assumptions; discussed practical implementation-side considerations; and outlined specific example targets for which this algorithm applies. The algorithm is flexible in the choice of $\Psi$, the space of $Y$, as well as the type of design.
@@ 4.1 Data Generation
    Quantify gains in statistical accuracy and efficiency as the sample size increases and compare SDRF with DRF [21], an existing distributional random forest that does not incorporate survey weights in the estimation process.
@@
$\bigl(\mathds{1}\{\mathrm{Strat}_{i}=1\},\ldots,\mathds{1}\{\mathrm{Strat}_{i}=H\}\bigr)\sim\operatorname{Multinomial}\!\left(1;\tfrac{1}{H},\ldots,\tfrac{1}{H}\right).$
@@
$M_{hj}\;=\;8\mathds{1}(Z_{hj}>0)+2\mathds{1}(\bar{Z}_{hj}\leq 0),\quad\text{where }\bar{Z}_{hj}\;=\;\frac{1}{N_{hj}}\sum_{i\in\mathrm{PSU}_{hj}}Z_{i}$
the PSU-level mean of $Z_{i}$, $N_{hj}=|\mathrm{PSU}_{hj}|$ the number of subjects within that PSU.

### v2

@@ 3.2 Splitting Rule
Fix a node region $P_{a}\subseteq\mathcal{X}$ and a candidate split $\theta=(j,t)$ that induces children $C_{L},C_{R}$. Define ${N}_{P_{a}}:=|\{i\in[N]:X_{i}\in P_{a}\}|$, $N_{\ell}:=|\{i\in[N]:X_{i}\in C_{\ell}\}|$. Suppose there exists $\alpha\in(0,1/2]$ such that $\frac{N_{L}}{N_{P_{a}}},\ \frac{N_{R}}{N_{P_{a}}}\in[\alpha,1-\alpha].$ Let $M^{N}(\theta)$ be the finite-population split score, $M_{n_{s}}(\theta)$ be the design-weighted sample score, and $M_{n_{s}}^{*(M)}(\theta)$ be the score using averaged bootstrap multipliers $\bar{n}_{i}^{*}:=M^{-1}\sum_{b=1}^{M}n_{b,i}^{*}$. Under (D1)–(D4), (R1)–(R3), and (K1)–(K2), if the multiplier increments $\delta_{b,i}:=n_{b,i}^{*}-1$ are conditionally sub-gamma and negatively dependent given $\xi$ (satisfied by the re-sampling methods described in Section S5.1, Supplementary Material, by Lemma S6, Supplementary Material), then
$\bigl|M_{n_{s}}^{*(M)}(\theta)-M^{N}(\theta)\bigr|=O_{p_{d}}\!\left(N_{P_{a}}^{-1/2}\;+\;\frac{1}{\sqrt{M}}\sum_{\ell\in\{L,R\}}n_{\mathrm{eff},\ell}^{-1/2}\right),$
where $n_{\mathrm{eff},\ell}=\frac{(\sum_{i\in S\bigcap C_{\ell}}\pi_{i}^{-1})^{2}}{\sum_{i\in S\bigcap C_{\ell}}\pi_{i}^{-2}}$ is the design-weighted effective sample size (Kish1995-aq) in child $\ell$.
@@
Proposition 3.3 separates two sources of split-score fluctuation: an intrinsic design term $|M_{n_{s}}(\theta)-M^{N}(\theta)|=O_{p_{d}}(N_{P_{a}}^{-1/2})$ which concerns the sampling noise already present even with fixed $n^{*}$, and a resampling term $|M_{n_{s}}^{*(M)}(\theta)-M_{n_{s}}(\theta)|=O_{p_{n^{*}|\xi}}\!\big(M^{-1/2}\sum_{\ell\in\{L,R\}}n_{\mathrm{eff},\ell}^{-1/2}\big)$.
The cross-over threshold occurs when these terms balance: if $M\gtrsim\tilde{M}$ where $\tilde{M}\asymp N_{P_{a}}\!\big(\sum_{\ell\in\{L,R\}}\,n_{\mathrm{eff},\ell}^{-1/2}\big)^{2},$ then the design term dominates and increasing $M$ further yields negligible stability gains. At deep nodes (small $n_{\mathrm{eff},\ell}$), this threshold is reached quickly: for instance, if both children have $n_{\mathrm{eff},\ell}\approx 100$ and $N_{P_{a}}\approx 200$, then $\tilde{M}\approx 8$ suffices. Beyond this point, split stability is governed by the effective sample sizes, controlled via minimum-node-size constraints, shallower depths, or larger subsampling fractions $q$, rather than by the number of resamples.
@@ 3.3 Prediction and Aggregation
    There exist constants $0<m<M<\infty$ such that for all radius r, of some $L(\mathbf{x})$ for the given $\mathbf{x}$ such that $0<r\leq k_{N}$, $mr^{p}\leq\#\{i:X_{i}\in B_{r}(\mathbf{x})\}/N\leq Mr^{p}$, where $B_{r}(\mathbf{x})$ is the ball centered at $\mathbf{x}$ with radius $r$.
@@
    For any $L^{*}$ found by $D_{split}$, when populating the values with $D_{est}$, all leaves receive at least $n_{L^{*},min}$ subjects. In other words, $n_{L^{*}(\mathbf{x})}=\sum_{i}^{N}\xi_{i}\mathbbm{1}(X_{i}\in L^{*}(\mathbf{x}))\geq n_{L^{*},min}>0$.
@@
    For any $\mathbf{x}$ and some deterministic chosen sequence $k_{N}\to 0$ such that $k^{p}_{N}N\to\infty$ as $N\to\infty$, there exists $t_{N}\to 0$ such that $P_{L}(diam(L(\mathbf{x}))\geq k_{N})\leq t_{N}$, where $diam(L(\mathbf{x}))=\sup_{x_{1},x_{2}\in L(\mathbf{x})}d(x_{1},x_{2})$, and $d$ is some metric on $\mathcal{X}^{p}$.
@@
    $\sup_{x\in\mathcal{X}_{1}\cup\mathcal{X}_{2}}\mathbb{P}\{X\in L^{*}(\mathbf{x})\,\Delta\,L(\mathbf{x})\}=o_{p}(d_{N})$ for some $d_{N}=o(k_{N}^{p})$, where $\Delta$ denotes symmetric difference: $A\Delta B:=(A\setminus B)\;\cup\;(B\setminus A)$.
Denote $\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ as the SDRF estimate. Then $d_{\mathbf{k}}(\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}),{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}))\to_{p_{d}}0$ as $N,B\to\infty$.
@@
Regularization–type conditions (A1)–(A3) can be controlled via adaptive stopping criteria. (A1) imposes a geometric control on the shape of $L$ to control the stretching. (A2) requires a local mass control for the local empirical density to be well-behaved. (A3) ensures that the local density estimate stays bounded. (B1)-(B3) formalize how the partitions produced by the algorithm behave as the finite population size grows. (B1) is a shrinking diameter assumption that mirrors the adaptive concentration result proved for i.i.d. forests Wager2015-nc. (B2) requires that the thickness of the error band be negligible relative to the leaf’s own diameter. In other words, the rate at which the leaf $L^{*}(\cdot)$ function estimated from the bootstrap sample approaches the leaf $L(\cdot)$ estimated when using the full finite population must be faster than the oracle leaf’s own shrinking-volume rate. Under the required relationship between $d_{N}$ and $k_{N}$, together with (A1) regularity conditions on the geometry, the symmetric-difference region $L^{*}(\mathbf{x})\triangle L(\mathbf{x})$ has measure $O_{p}(d_{N}/k_{N}^{p})$, which implies that both the tree estimator $\widehat{T}^{*}_{N}(L_{\xi}^{*}(\mathbf{x}))$ and its conditional variance coincide with their oracle counterparts up to $o_{p}(N^{-1/2})$ and $o_{p}(N^{-1})$, respectively. Among all the assumptions, (B2) is the hardest to directly control. Thus we have provided Theorem 3.2 that partially shed light on conditions under which (B2) can be reasonably assumed to hold.
@@
Suppose $\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{N}$ are i.i.d. realizations from $P_{\Omega}(\cdot)$. Suppose $\mathbf{k}$ satisfies (K1)–(K3), and $\mathcal{Y}$ is locally compact Polish (satisfying (S1)). Denote ${P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ as the true conditional measure of $Y|X=\mathbf{x}$ in the design space and $\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ is an estimate for it. Denote $P_{\Omega}(\mathbf{x})$ as the true conditional measure of $Y|X=\mathbf{x}$ in the super population space. Then for any specified $\mathbf{x}\in\mathcal{X}$ and $\Psi:(\mathcal{M}_{b}(\mathcal{Y}),\;d_{\mathbf{k}})\to(\mathcal{V},d)$ a measurable functional continuous at $P^{N}(Y|X=\mathbf{x})$ for a.e. $\omega\in\Omega$, if   $d_{\mathbf{k}}(\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}),{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}))\to_{p_{d}}0,\text{ then }$
@@
Corollary 3.5 reduces design– and model–based consistency of $\widehat{\Psi}_{N}(\mathbf{x})$ to control of $\widehat{P}_{\mathcal{S}^{N}\mid\omega}(\mathbf{x})$ in the $d_{\mathbf{k}}$-metric, together with continuity of the functional $\Psi$ at the relevant target law. Many standard summaries satisfy this requirement under mild regularity. For example, when $\mathcal{Y}=\mathbb{R}^{d}$, linear functionals such as conditional means $\Psi(P)=\int y\,dP(y)$ and, more generally, $\Psi_{g}(P)=\int g(y)\,dP(y)$ for $g\in C_{b}(\mathcal{Y})$ are continuous under weak convergence. For distributional functionals, $\Psi_{t}(P)=F_{P}(t)$ and $\Psi_{t}(P)=P(Y>t)$ are continuous at $P$ whenever $P(\{t\})=0$. The quantile functional $\Psi_{q}(P)=F_{P}^{-1}(q)$ is continuous at $P$ whenever the $q$-quantile is unique. Equivalently, that $F_{P}$ has no flat segment at level $q$. The kernel mean element itself and its norm $\Psi(P)=\|\mu_{\mathbf{k}}(P)\|_{\mathcal{H}}$, and the MMD between the true conditional $P$ to a reference distribution $Q$, $d_{\mathbf{k}}(P,Q)$, are also examples. For higher-order continuous functionals such as covariances or variances of kernel embeddings, larger effective sample sizes may be needed to ensure adequate concentration of the estimator $\widehat{P}_{\mathcal{S}^{N}|\omega}$ in the $d_{\mathbf{k}}$-metric.
@@ 3.4 Implementation Considerations
  • For the re–sampling, Conti2020-mc showed that asymptotically, generating one pseudo population and then drawing $B$ samples from it (equivalently, realizing $B$ times $n^{*}$ conditioning on the realized pseudo population) is asymptotically equivalent to realizing independently $B$ times both the pseudo population “and” the re-sample.
  • RBF–type kernel is a common class of kernels that is translation–invariant. The median-heuristic, i.e., specifying the bandwidth $\sigma>0$ to be the median of pairwise sampled distance, is a commonly adopted, and reasonable choice (Garreau2017-ro; matabuena2023hypothesis).
  • Hyperparameters governing the stopping criteria are introduced to help ensure regularity-type conditions (A1)–(A3). For finite samples, it may be preferable to use a moderate number of shallow trees to keep local sample weights well-controlled. Moreover, when the number of PSUs is large but PSU sizes are small, one may consider omitting second-stage resampling (Section S5.1, Supplementary Material). We note that the maximum node–level weight ratio $\max_{i\in\mathcal{I}(L)}\hat{w}_{i}/\min_{i\in\mathcal{I}(L)}\hat{w}_{i}$ where $\mathcal{I}(L)$ the subject index set for node $L$ is the only additional hyperparameter compared to the classical i.i.d. random forests for keeping the weights controlled.
  • Because SDRF yields a kernel-type estimator, one could further post-process $\widehat{P}(Y\mid X=\mathbf{x})$ (or $\widehat{\Psi}(\mathbf{x})$) with an additional kernel-smoothing step verdinelli2021forest in $\mathbf{x}$ to reduce finite-sample variability and stabilize the resulting estimates, particularly when it is reasonable to assume that the underlying target varies smoothly in $\mathbf{x}$.
To conclude, up to this point, we have illustrated the necessity of an appropriate method when data arise under sampling designs as opposed to being i.i.d., specifically when $Y\not\!\perp\!\!\!\perp Z\mid X$ and $\xi\not\!\perp\!\!\!\perp Z\mid X$ (Section S2, Supplementary Material); described in detail our proposed algorithm as a design-aware solution; proved both types of consistency for the resulting estimator under reasonable assumptions; discussed practical implementation-side considerations; and outlined specific example targets for which this algorithm applies. The algorithm is flexible in the choice of $\Psi$, the space of $Y$, as well as the type of design.
@@ 4.1 Data Generation
    Quantify gains in statistical accuracy and efficiency as the sample size increases and compare SDRF with DRF cevid_michel_meinshausen_bühlmann_2022, an existing distributional random forest that does not incorporate survey weights in the estimation process.
@@
$\bigl(\mathbbm{1}\{\mathrm{Strat}_{i}=1\},\ldots,\mathbbm{1}\{\mathrm{Strat}_{i}=H\}\bigr)\sim\operatorname{Multinomial}\!\left(1;\tfrac{1}{H},\ldots,\tfrac{1}{H}\right).$
@@
$M_{hj}\;=\;8\mathbbm{1}(Z_{hj}>0)+2\mathbbm{1}(\bar{Z}_{hj}\leq 0),\quad\text{where }\bar{Z}_{hj}\;=\;\frac{1}{N_{hj}}\sum_{i\in\mathrm{PSU}_{hj}}Z_{i}$
is the PSU-level mean of $Z_{i}$, $N_{hj}=|\mathrm{PSU}_{hj}|$ the number of subjects within that PSU.